# Analyzing and Mitigating Bias for Vulnerable Classes: Towards Balanced Representation in Dataset

### v1

@@ Abstract
The accuracy and fairness of perception systems in autonomous driving are crucial, particularly for vulnerable road users. Mainstream research has looked into improving the performance metrics for classification accuracy. However, the hidden traits of bias inheritance in the AI models, class imbalances and disparities in the datasets are often overlooked. In this context, our study examines the class imbalances for vulnerable road users by focusing on class distribution analysis, performance evaluation, and bias impact assessment. We identify the concern of imbalances in class representation, leading to potential biases in detection accuracy. Utilizing popular CNN models and Vision Transformers (ViTs) with the nuScenes dataset, our performance evaluation reveals detection disparities for underrepresented classes. We propose a methodology for model optimization and bias mitigation, which includes data augmentation, resampling, and metric-specific learning. Using the proposed mitigation approaches, we see improvement in IoU(%) and NDS(%) metrics from 71.3 to 75.6 and 80.6 to 83.7 respectively, for the CNN model. Similarly, for ViT, we observe improvement in IoU and NDS metrics from 74.9 to 79.2 and 83.8 to 87.1 respectively. This research contributes to developing more reliable models and datasets, enhancing inclusiveness for minority classes.
@@ I Introduction
The advent of autonomous driving technology marks a significant leap forward in the automotive industry. While the benefits of autonomous vehicles (AVs) are manifold, including potential reductions in traffic accidents and increased mobility for various demographics, the technology’s reliability is highly contingent on the accuracy and fairness of its perception systems [12] [14]. These systems, particularly those involving object detection, are critical for identifying and appropriately reacting to various entities on the road, such as pedestrians, cyclists, motorcyclists, and mobility-impaired persons [8] [31]. However, recent studies have underscored the existence of biases in autonomous driving datasets, which could lead to discriminatory treatment of certain road users [17] [25]. Our study delves into two prominent datasets in this domain, nuScenes [4] and Waymo [27], aiming to uncover and address these biases. We explore the representation of vulnerable classes on object detection models’ accuracy and fairness through class distribution analysis, model’s performance evaluation, and bias impact assessment. Class imbalance refers to the disproportionate representation of different categories or classes within a dataset [23]. In the context of autonomous driving, this often manifests as an overrepresentation of certain types of vehicles, while pedestrians, cyclists, motorcyclists, and especially mobility-impaired individuals are underrepresented [17]. The imbalance is not only a statistical concern but a practical one, as it can lead to perception systems that are less adept at recognizing minor, yet crucial classes. Studies have shown that this imbalance can result in biased learning outcomes, where models preferentially detect overrepresented classes [29]. The consequences of such an imbalance in the dataset cannot be overlooked, as evidenced by incidents like the Uber self-driving car collision [26]. In this dataset imbalance context, our paper explores:
  1. 1.
@@
To counter biases towards learning, our research explores a framework for balanced class representation and AI model optimization. We emphasize the need for diverse and representative datasets that include a wide range of scenarios, reflecting the complexity of real-world conditions. The proposed framework includes bias assessment and model re-calibration to ensure model fairness and effectiveness with the retraining of models with the inclusion of new classes/sub-samples. We also highlight critical gaps in current vision datasets used in autonomous driving through metric-specific learning to develop more equitable and reliable systems. The aim is to contribute to the generalization of AI algorithms and respective datasets.
@@ II Background and Related Work
Perception technology in autonomous driving has evolved through extensive research and development, with significant milestones achieved in sensor technology, algorithmic processing, and machine learning models. Early systems relied heavily on rule-based algorithms and limited sensor input, but the advancements in deep learning and improvements in LiDAR, radar, and camera technologies have provided next-generation solutions [15]. Despite these advancements, the challenge of developing perception systems that can reliably interpret diverse and unpredictable road scenarios remains a significant research focus. Datasets play a crucial role in training and validating autonomous driving systems. Datasets like KITTI [9] and Cityscapes [6] have provided foundations for advancing the field. More recently, datasets such as nuScenes and Waymo have offered more diverse and complex environments for training and testing AI models. These datasets include various annotations for vehicles, pedestrians, cyclists, and other entities, providing a rich foundation for developing and evaluating perception algorithms.
@@
Class Imbalance: Efforts to address class imbalance and bias in autonomous driving datasets have included techniques like data augmentation, re-sampling, and advanced algorithmic approaches like cost-sensitive learning [11]. Research in this area seeks to balance the representation of classes in training data [28] and develop fair and accurate models across all categories [29] [31]. For instance, in research paper[10], the author highlights the use of scenario simulations to test and improve model performance in detecting imbalanced classes. The ethical implications of biased autonomous driving systems are significant. There is a growing discourse on the need for equitable AI systems that ensure the safety and fairness of all road users [24]. Furthermore, regulatory bodies are increasingly focusing on how these technologies adhere to safety standards, particularly in terms of their ability to detect and respond to diverse and representative [5] scenarios. Targeting imbalanced representation, a novel approach that combines visual codebook generation with deep features and a non-linear $Chi^{2}$ SVM classifier is porposed in [23]. This method tackles the issue of imbalanced datasets, where traditional algorithms often fail to detect minority classes effectively. The approach involves extracting low-level deep features using transfer learning with the ResNet-50 pre-trained network and k-means clustering to create a visual codebook. Each image is then represented as a Bag-of-Visual-Words (BOVW), derived from the histogram of visual words in the vocabulary. The $Chi^{2}$ SVM classifier is used for classifying these features, demonstrating optimal performance in empirical analysis. This method shows superior accuracy, F1-score, and AUC metrics results compared to state-of-the-art methods, as validated on two challenging datasets: Graz-02 and TF-Flowers [23].
@@
With a focus on training methods to address the challenges of learning from imbalanced data, a new loss function that mitigates the impact of samples leading to overfitted decision boundaries is proposed in [21]. This loss function has been shown to enhance the performance of various imbalance learning methods. The proposed approach is versatile and can be integrated with existing resampling, meta-learning, and cost-sensitive learning methods to tackle class imbalance problems. The paper validates this method with experiments on multiple benchmark datasets, demonstrating its superiority over state-of-the-art cost-sensitive loss methods. Additionally, the authors have made their code publicly available, facilitating further research and application [21]. A novel approach for predicting driver behaviour, which is crucial for safely integrating autonomous vehicles into human-dominated traffic, is presented in [3]. The proposed method addresses the shortcomings of existing predictive models, which either lack transparency (deep neural networks) or are not expressive enough (rule-based models). The authors introduce a model that embeds the Intelligent Driver Model (IDM), a rule-based approach, into deep neural networks. This hybrid model combines the long-term coherence and interpretability of rule-based models with the expressiveness of deep learning, aiming to accurately predict driver behaviour in complex scenarios like merging. The method is an attempt to bridge the gap between two modelling paradigms. It enhances the interpretability of neural network predictions while maintaining accuracy, a critical factor for real-time decision-making in autonomous driving. The model’s transparency is particularly beneficial for debugging and understanding its predictions.
@@ III Methodology
Figure 1: Proposed methodology for analyzing and mitigating biases
@@ III-A Dataset and Classes
In this paper, we use nuScenes as the case study, focusing on the Pedestrian, Cyclist, and Motorcyclist classes. However, the approach can be tested and validated on other driving datasets, e.g., the Waymo dataset [27], where the samples can be Vehicles, Pedestrians, and Cyclists.
III-A1 3.1.1 Class Distribution Analysis
We perform statistical analysis on the frequency of the specified classes in the nuScenes and Waymo datasets to identify class imbalances. The datasets are scrutinized to understand the context (urban versus rural settings, diverse weather and lighting conditions) in which these classes are represented, highlighting any underrepresented scenarios.
Pedestrian Class: nuScenes dataset includes a significant proportion of pedestrian annotations, with 149,921 instances labelled as “human.pedestrian.adult” accounting for 21.61% of all annotations. This substantial representation highlights the importance of pedestrian detection in autonomous driving systems. However, other pedestrian sub-categories like children (1,934 annotations, 0.28%), construction workers (13,582 annotations, 1.96%), and individuals with personal mobility devices (2,281 annotations, 0.33%) are less represented. These differences underscore potential gaps in the dataset’s and subsamples diversity.
Cyclist Class: It is categorized as “vehicle.bicycle” and has around 17,060 annotations, comprising 2.46% of the total dataset. This relatively lower representation compared to pedestrians impacts the class weight calculation and therefore the model’s ability to accurately detect cyclists, which might become a critical safety concern in urban driving scenarios.
@@
The common metrics used for model performance evaluation in nuScenes are, mean average Precision (mAP), Intersection over Union (IoU), and nuScenes Detection Score (NDS). mAP and IoU are well-known metrics used in 3d object detection and segmentation tasks [4], and are calculated using the fundamentals of precision, recall, area of overlap and intersection over union respectively. Similarly, NDS combines several metrics, including the mAP, to provide a comprehensive performance score for the object detection models. The NDS metric is calculated as:
@@
III-A2 3.1.2 Data Augmentation/Re-sampling
Given the class representation imbalances, particularly in the less represented sub-categories of pedestrians and the overall lower representation of cyclists and motorcyclists, specific strategies can be employed for model training and evaluation.
Resampling Technique: Addressing class imbalances in the datasets, particularly for underrepresented classes like cyclists and motorcyclists, is crucial for unbiased model training in autonomous driving systems. We apply both “Random Oversampling” and “Undersampling ” from resampling techniques”. Random Oversampling duplicates instances from minority classes, ensuring models have sufficient data to learn from these critical yet less frequent road users. Conversely, Undersampling reduces the predominance of over-represented classes, like pedestrians, to prevent bias inheritance in models. Both methods are carefully balanced to maintain data diversity [20].
Data Augmentation: Another key approach is the use of geometric transformations, such as rotating and flipping images. This technique represents various class orientations, which is crucial for autonomous driving where objects are encountered from multiple directions. These transformations not only prevent the model from developing orientation-specific biases but also enhance its ability to adapt to diverse real-world scenarios. Additionally, this approach contributes to increasing the dataset’s diversity and size, thereby improving the model’s robustness and generalization capabilities for more accurate and reliable perception tasks [22].
Combination of Strategies: To further have varied training and testing and to increase the samples/sub-classes diversity, we combine resampling and data augmentation to balance the training dataset. This approach integrates the strengths of both Resampling (Random undersampling/Oversampling) and Transformations, such as rotating and flipping images, to optimize the representation and variability of data.
@@ III-C Bias Impact Assessment
This section describes model errors (false positives and negatives) and examines learning patterns and bias inheritance, crucial for identifying potential biases in model performance.
III-C1 3.3.1 Error Analysis
We focus on false positives (e.g., incorrectly identifying objects as belonging to a target class) and false negatives (e.g., failing to detect cyclists). We further use Class-specific error analysis, which includes evaluating how well the model detects pedestrians, cyclists, and motorcyclists under various scenarios. We also examine if there is a tendency to overlook certain classes and how this impacts overall model performance. The correlation between these errors and previously identified biases, such as neurons’ sensitivity or selectivity scores, is also explored to understand the underlying causes and guide bias mitigation strategies using bias correlation.
III-C2 3.3.2 Learning Patterns and Bias Inheritance
Different model architectures lead to varied learning behaviours and potential biases. CNNs like ResNet may focus on textures and local patterns, which could introduce biases if such features are not uniformly present across all classes. This can result in misclassification, especially when distinguishing features are missing. In contrast, Vision Transformers (ViTs) may develop biases based on how attention is distributed across an image.
@@
Interpretation and Visualization: Layer-wise Relevance Propagation (LRP) provides insights into what drives a model’s decisions by highlighting influential input features. For example, if a model consistently focuses on irrelevant background features for decision-making, it indicates a bias that needs correction [19]. Analysis of these scores, using visual tools like heatmaps and attention maps, helps in interpreting these complex models, offering an understanding of their decision-making processes, and adjusting the training process to mitigate identified biases.
Overall, this process involving analysis of learning patterns, data bias inheritance comparative model analysis, and visualization can help discover bias occurrence/inheritance in different model architectures.
@@ III-D Bias Mitigation Techniques
To mitigate biases inherited from sample representation, we explore the strategies mentioned below to address these class disparities. These techniques are crucial for addressing the disparities, ensuring that the model’s performance is equitable and reliable across all categories.
III-D1 3.4.1 Cost-Sensitive Learning
Customized cost-sensitive learning offers a robust approach to mitigating these issues by adjusting the model’s loss function according to the representation of each class. The class weights calculation and loss function are discussed below.
@@
By employing this customized cost-sensitive learning approach, we aim to reduce the biases inherent in the AI model from the imbalanced dataset, ensuring a more equitable and robust model performance across all classes, which is crucial for the reliability and safety of perception applications.
III-D2 3.4.2 Data Augmentation Based on Model Analysis
We use a targeted data augmentation strategy to address biases identified in ViTs through attention map analysis and Layer-wise Relevance Propagation (LRP). This approach specifically addresses under-representation issues in cyclists and motorcyclists within the nuScenes dataset.
@@
LRP-Informed Sampling: Layer-wise Relevance Propagation provides a deeper understanding of which features contribute most significantly to the model’s decisions. Using insights from LRP, we can selectively augment our dataset to enhance the representation of features critical for correct classifications. Augmentation for this case includes scenarios where the model typically misclassifies images, like partially occluded subjects or specific textures and patterns, enhancing the model’s ability to differentiate between relevant and misleading features.
Implementation:This strategy involves iterative dataset refinement based on continuous model analysis. By aligning augmentation closely with model needs and performance, we aim to mitigate biases, effectively ensuring balanced and fair model performance. This is critical for perception systems, where the accurate detection of all road users, particularly those underrepresented like cyclists and motorcyclists.

### v2

@@ Abstract
The accuracy and fairness of perception systems in autonomous driving are essential, especially for vulnerable road users such as cyclists, pedestrians, and motorcyclists who face significant risks in urban driving environments. While mainstream research primarily enhances class performance metrics, the hidden traits of bias inheritance in the AI models, class imbalances and disparities within the datasets are often overlooked. Our research addresses these issues by investigating class imbalances among vulnerable road users, with a focus on analyzing class distribution, evaluating performance, and assessing bias impact. Utilizing popular CNN models and Vision Transformers (ViTs) with the nuScenes dataset, our performance evaluation indicates detection disparities for underrepresented classes. Compared to related work, we focus on metric-specific and Cost-Sensitive learning for model optimization and bias mitigation, which includes data augmentation and resampling. Using the proposed mitigation approaches, we see improvement in IoU(%) and NDS(%) metrics from 71.3 to 75.6 and 80.6 to 83.7 for the CNN model. Similarly, for ViT, we observe improvement in IoU and NDS metrics from 74.9 to 79.2 and 83.8 to 87.1. This research contributes to developing reliable models while enhancing inclusiveness for minority classes in datasets.
@@ I Introduction
Autonomous driving is dependent on several sensory and AI technologies. The benefits of deploying autonomous vehicles (AVs) have been associated with traffic optimization and safe manoeuvre. However, the fundamentals of such technologies highly rely on the accuracy and fairness of their perception systems [13, 16]. These systems, particularly those involving object detection, are critical for identifying various objects or classes on the road, such as pedestrians, cyclists, motorcyclists, and mobility-impaired persons [9, 34, 15]. Recent studies have highlighted the existence of biases in vision datasets, including autonomous driving datasets, which could lead to discriminatory treatment of certain road users [19, 28]. Within this focus, our study discusses two prominent datasets in this domain, nuScenes [5] and Waymo [30], aiming to address class imbalances and possibilities of biases inheritance in a model when trained with class disparities. This paper covers the representation of vulnerable classes in object detection models’ accuracy and fairness through class distribution analysis, model performance evaluation, and bias impact assessment. Class imbalance refers to the disproportionate representation of different categories or classes within a dataset [26, 4]. Within the context, this often manifests as an over-representation of type of class such as vehicles, while pedestrians, cyclists, motorcyclists, and especially mobility-impaired individuals are classes that can remain underrepresented [19]. This imbalance can lead to perception systems that are less adept at recognizing minor yet crucial classes [21, 19, 2]. Studies have shown that this imbalance can result in the inheritance of bias in AI models from the datasets, where models preferentially detect overrepresented classes [32]. Figure 1 shows a heatmap from our tests, where images on the left are input images and gradients flow for a class is shown on the right. The consequences of such an imbalance in the dataset cannot be overlooked, as evidenced by incidents like the Uber self-driving car collision [29]. We emphasize the need for balanced, diverse and representative datasets that include a wide range of scenarios, reflecting the complexity of real-world conditions. In this dataset imbalance context, our paper explores:
@@
Figure 1: Three samples that show testing results when a single class is predominantly present during model training.
Compared to previous works, which are dependent on the analysis of “F-1 Score, False Positives(FP) and False Negatives(FN)” for bias impact assessment tasks, our research explores sensitivity and selectivity Score for NNs and Layer-wise relevance propagation for vision transformer models within a framework for class-specific learning/evaluation process to identify bias inheritance. The framework then includes bias assessment using FP, FN, and model re-calibration process to ensure model fairness and effectiveness if a model is retrained on new classes/sub-samples. We also highlight critical gaps in current vision datasets used in autonomous driving through metric-specific learning to develop more equitable and reliable systems. The aim is to contribute to generalising AI algorithms and respective datasets.
@@ II Background and Related Work
Perception technology in autonomous driving has evolved through extensive research and development, with milestones achieved in sensor technology, algorithmic processing, and machine learning models. Early systems relied heavily on rule-based algorithms and limited sensor input, but the advancements in deep learning and improvements in LiDAR, radar, and camera technologies have provided next-generation solutions [17]. Despite these advancements, the challenge of developing perception systems that can reliably interpret diverse and unpredictable road scenarios remains an important focus. Datasets play a crucial role in training and validating autonomous driving systems. Datasets like KITTI [10] and Cityscapes [7] have provided foundations for advancing the field. In recent years, datasets such as nuScenes and Waymo have offered more diverse and complex model training and testing environments. These datasets include various annotations for vehicles, pedestrians, cyclists, and other entities, providing a rich foundation for developing and evaluating perception algorithms.
@@
Class Imbalance: Efforts to address class imbalance and bias in autonomous driving datasets have included techniques like data augmentation, re-sampling, and advanced algorithmic approaches like cost-sensitive learning [12, 4, 21]. Research in this area seeks to balance the representation of classes in training data [31] and develop fair and accurate models across all categories [32] [34]. In paper[11], the author highlights the use of scenario simulations to test and improve model performance in detecting imbalanced classes. The ethical implications of biased autonomous driving systems are significant. There is a growing discourse on the need for equitable AI systems that ensure the safety and fairness of all road users [27]. Furthermore, regulatory bodies are increasingly focusing on how these technologies adhere to safety standards, such as their ability to detect and respond to diverse and representative [6] scenarios. Targeting imbalanced representation, an approach that combines visual codebook generation with deep features and a non-linear $Chi^{2}$ SVM classifier, is proposed in [26]. This method tackles the issue of imbalanced datasets, where algorithms often fail to detect minority classes. The approach extracts low-level deep features using transfer learning with the ResNet-50 pre-trained model and k-means clustering to create a visual codebook. Each image is then represented as a Bag-of-Visual-Words (BOVW), derived from the histogram of visual words in the vocabulary. The $Chi^{2}$ SVM classifier is used for classifying these features, demonstrating optimal performance in empirical analysis. This method shows superior accuracy, F1-score, and AUC metrics results compared to state-of-the-art methods, as validated on two challenging datasets: Graz-02 and TF-Flowers [26].
@@
With a focus on training methods to address the challenges of learning from imbalanced data, a new loss function that mitigates the impact of samples leading to overfitted decision boundaries is proposed in [24]. This loss function has been shown to enhance the performance of various imbalance learning methods. The proposed approach is versatile and can be integrated with existing resampling, meta-learning, and cost-sensitive learning methods to tackle class imbalance problems. The paper validates this method with experiments on multiple benchmark datasets, demonstrating its superiority over state-of-the-art cost-sensitive loss methods. Additionally, the authors have made their code publicly available, facilitating further research and application [24]. An approach for predicting driver behaviour, which is crucial for safely integrating autonomous vehicles into human-dominated traffic, is presented in [3]. The proposed method addresses the shortcomings of existing predictive models, which either lack transparency (deep neural networks) or are not expressive enough (rule-based models). The authors introduce a model that embeds the Intelligent Driver Model (IDM), a rule-based approach, into deep neural networks. This hybrid model combines the long-term coherence and interpretability of rule-based models with the expressiveness of deep learning, aiming to accurately predict driver behaviour in complex scenarios like merging. The method is an attempt to bridge the gap between two modelling paradigms. It enhances the interpretability of neural network predictions while maintaining accuracy, a critical factor for real-time decision-making in autonomous driving. The model’s transparency is particularly beneficial for debugging and understanding its predictions.
@@
Figure 2: Proposed methodology for analyzing and mitigating biases
@@ III-A Dataset and Classes
In this paper, we use nuScenes as the case study, focusing on the Pedestrian, Cyclist, and Motorcyclist classes. However, the approach can be tested and validated on other driving datasets, e.g., the Waymo dataset [30], where the samples and classes can be varied.
A) Class Distribution Analysis: We perform statistical analysis on the frequency of the specified classes in the nuScenes and Waymo datasets to identify class imbalances. The datasets are scrutinized to understand the context (urban versus rural settings, diverse weather and lighting conditions) in which these classes are represented, highlighting any underrepresented scenarios.
Pedestrian Class: nuScenes dataset includes a major proportion of pedestrian annotations, with 149,921 instances labelled as “human.pedestrian.adult” accounting for 21.61% of all annotations. This substantial representation highlights the importance of pedestrian detection in autonomous driving systems. However, other pedestrian sub-categories like children (1,934 annotations, 0.28%), construction workers (13,582 annotations, 1.96%), and individuals with personal mobility devices (2,281 annotations, 0.33%) are less represented. These differences shows gaps in dataset and subsamples diversity.
Cyclist Class: It is categorized as “vehicle.bicycle” and has around 17,060 annotations, comprising 2.46% of the total dataset. This relatively lower representation compared to pedestrians impacts the class weight calculation and, therefore, the model’s ability to accurately detect cyclists, a safety concern in urban driving scenarios.
@@
The standard metrics used for model performance evaluation in nuScenes are mean average Precision (mAP), Intersection over Union (IoU), and nuScenes Detection Score (NDS). mAP and IoU are well-known metrics used in 3d object detection and segmentation tasks [5, 15], and are calculated using the fundamentals of precision, recall, area of overlap and intersection over union respectively. Similarly, NDS combines several metrics, including the mAP, to provide a comprehensive performance score for the object detection models. The NDS metric is calculated as:
@@
B) Data Augmentation/Re-sampling: Given the class representations, particularly for the less represented sub-categories of pedestrians and overall lower representation of cyclists and motorcyclists, following strategies can be used for model training and evaluation.
Resampling Technique: Addressing class imbalances in the datasets for underrepresented classes, like cyclists and motorcyclists, is crucial for unbiased training in autonomous driving models. We apply both “Random Oversampling” and “Undersampling ” from resampling techniques”. Random Oversampling duplicates instances from minority classes, ensuring models have sufficient data to learn from these critical yet less frequent road users. Conversely, Undersampling reduces the predominance of over-represented classes, like pedestrians, to prevent bias inheritance in models. Both methods are carefully balanced to maintain data diversity [23].
Data Augmentation: Another key approach is the use of geometric transformations, such as rotating and flipping images. This technique represents various class orientations, which is crucial for autonomous driving where objects are encountered from multiple directions. These transformations prevent the model from developing orientation-specific biases and enhance its ability to adapt to diverse real-world scenarios. Additionally, this approach contributes to increasing the dataset’s diversity and size, thereby improving the model’s robustness and generalization capabilities for more accurate and reliable perception tasks [25].
Combination of Strategies: For varied training, testing, and to increase the samples/sub-classes diversity, we combine resampling and data augmentation to balance the training dataset. This approach integrates the strengths of both Resampling (Random undersampling/Oversampling) and Transformations, such as rotating and flipping images, to optimize the representation and variability of data.
@@ III-C Bias Impact Assessment
This section describes model errors (false positives and negatives) and examines learning patterns and bias inheritance, which are crucial for identifying potential biases in model performance.
A) Error Analysis: We focus on false positives (e.g., incorrectly identifying objects as belonging to a target class) and false negatives (e.g., failing to detect cyclists). We further use Class-specific error analysis, which includes evaluating how well the model detects pedestrians, cyclists, and motorcyclists under various scenarios. We also examine if there is a tendency to overlook certain classes and how this impacts overall model performance. The correlation between these errors and previously identified biases, such as neurons’ sensitivity or selectivity scores, is also explored to understand the underlying causes and guide bias mitigation strategies using bias correlation.
B) Learning Patterns and Bias Inheritance: Different model architectures lead to varied learning behaviours and potential biases. CNNs like ResNet may focus on textures and local patterns, which could introduce biases if such features are not uniformly present across all classes. This can result in misclassification, especially when distinguishing features are missing. In contrast, Vision Transformers (ViTs) may develop biases based on how attention is distributed across an image.
@@
Interpretation and Visualization: Layer-wise Relevance Propagation (LRP) provides insights into what drives a model’s decisions by highlighting influential input features. For example, if a model consistently focuses on irrelevant background features for decision-making, it indicates a bias that needs correction [22]. Analysis of these scores, using visual tools like heatmaps and attention maps, helps in interpreting these complex models, offering an understanding of their decision-making processes, and adjusting the training process to mitigate identified biases. Overall, this process involves the analysis of learning patterns, data bias inheritance, comparative model analysis, and visualization, which can help discover bias occurrence/inheritance in different model architectures.
@@ III-D Bias Mitigation Techniques
We explore the strategies mentioned below to mitigate biases inherited from sample representation and address these class disparities. These techniques are necessary for addressing the disparities and ensuring that the model’s performance is equitable and reliable.
A) Cost-Sensitive Learning: Customized cost-sensitive learning offers a robust approach to mitigating these issues by adjusting the model’s loss function according to the representation of each class. The class weights calculation and loss function are discussed below.
@@
By using this customized cost-sensitive learning approach, we aim to reduce the biases inherent in the AI model from the imbalanced dataset, ensuring a more equitable and robust model performance across all classes, which is crucial for the reliability and safety of perception applications.
B) Data Augmentation Based on Model Analysis: We use a targeted data augmentation strategy to address biases identified in ViTs through attention map analysis and Layer-wise Relevance Propagation (LRP). This approach specifically addresses under-representation issues in cyclists and motorcyclists within the nuScenes dataset.
@@
LRP-Informed Sampling: Layer-wise Relevance Propagation provides an understanding of which features contribute most to the model’s decisions. Using insights from LRP, the dataset can be augmented to enhance the representation of features critical for correct classifications. Augmentation for this case includes scenarios where the model typically misclassifies images, like partially occluded subjects or specific textures and patterns, enhancing the model’s ability to differentiate between relevant and misleading features.
Implementation: This strategy involves iterative dataset refinement based on continuous model analysis. By aligning augmentation closely with model performance, we aim to mitigate biases, ensuring balanced and fair model performance. This is critical for perception systems, where the accurate detection of all road users, particularly those underrepresented like cyclists and motorcyclists.